Need all parts some hints are giving ! If you don’t know don’t answer need experts only
Question # 5 (A) Two identical routers are located at site-A with 8 ports and at site-B with 10 ports. In how many ways can these routers be connected to 5 access points (Aps) placed on the network at the same time? Answer-hints: A and Bare independent. Jointly, they can enable 14,112 connections] (B) A Cisco edge-router in a WAN receives requests for connectivity on an average of 6 per minute. Assume each request is independent of other requests. Determine the probability of getting exactly 8 requests in one minute. Answer-hints: Poisson process;A-6 calls per minute; p(x 8) 0.103]
Answer #1

ques 5

a)

at site A, with 8 ports

no. of ways by which 5 access points can be connected=C(8,5)=56

at site B, with 10 ports

no. of ways by which 5 access points can be connected=C(10,5)=252

since,A and B are independent

so , n(A and B) = n(A)*n(B)

so, total number of ways by which these routers can be connected to 5 acccess points = 56*252 = 14112

------------------------------------

b)

it is poisson probability distribution.

in this ,

P(X=x) = eλx/x!

Mean/Expected number of requests, λ=    6 per minute

P ( X =    8   ) = e ^( -6) * 6 ^ 8 /8! = 0.1033